the length of a rug is twice its width. its area is 12.5 sq. meters. find the dimensions of the rug.

Answer by amarjeeth123(569)

Let the width of the rug be x.
Length of the rug is 2x.
Area of the rug=2x*x=2x^2
2x^2=12.5
x^2=6.25
x=2.5
The length of the rug is 5 meters and the width of the rug is 2.5 meters.